OpenStudy (anonymous):

A roller in printing press, with a shape of a solid cylinder and has a mass of 5 kg, turns through an angle θ(t) given by θ(t) = γt^(2) – βt^(3), where γ = 3.20 rad/s^(2) and β = 0.500 rad/s^(3). a. What is the maximum positive angular velocity and b. at what value of t does it occur?

OpenStudy (anonymous):

is it correct? a. angular velocity = 20.48 rad/s b. t = 6.4 s

OpenStudy (irishboy123):

are you using calculus?!?! because from \( θ(t) = γt^2 – βt^3\) , we have \(\dot θ(t) = 2γt – 3βt^2\) and \(\ddot θ(t) = 2γ – 6βt = 0 \implies t_{\dot \theta_{max}} = \dfrac{\gamma}{3 \beta} = 2.1\dot 3 \, s\) for max ang vel. [\(\dddot {\theta} <0 \) so it is a max] have a look here at plot of ang velocity (v) and ang acceleration (a) to clarify.... i was wondering how you got a) before b) and suspect maybe you are plotting it ....or maybe i am just really missing the point :-)) lol

OpenStudy (irishboy123):

this is the plot i referred to but failed to include: 

OpenStudy (irishboy123):

😋